OpenStudy (anonymous):

simplify the expression. 6a^10b^-9 ___________ 3a^4b-3 Write your answer without negative exponents.

OpenStudy (anonymous):

helpppp plz

OpenStudy (sidsiddhartha):

\[\frac{ 6a^{10}b^{-9} }{ 3a^4b^{-3} }=2*a^{10-4}*b^{-9+3}=2a^6b^{-6}=\frac{ 2a^6 }{ b^6 }\]

OpenStudy (sidsiddhartha):

@stacesweet

OpenStudy (sidsiddhartha):

used these properties\[\frac{ x^a }{ x^b }=x^{a-b}\\ \frac{ 1 }{ x^a }=x^{-a}\]